# Data-driven computation of adjoint sensitivities without adjoint solvers: An application to thermoacoustics

### v1

@@ Abstract
Adjoint methods have been the pillar of gradient-based optimization for decades. They enable the accurate computation of a gradient (sensitivity) of a quantity of interest with respect to all system’s parameters in one calculation. When the gradient is embedded in an optimization routine, the quantity of interest can be optimized for the system to have the desired behaviour. Adjoint methods, however, require the system’s Jacobian, whose computation can be cumbersome, and is problem dependent. In this paper, we propose a computational strategy to infer the adjoint sensitivities from data (observables), which bypasses the need of the Jacobian of the physical system. The key component of this strategy is an echo state network, which learns the dynamics of nonlinear regimes with varying parameters, and it evolves dynamically via a hidden state. Although the framework does not make assumptions on the dynamical system, we focus on thermoacoustics, which are governed by nonlinear and time-delayed systems. First, we show that a parameter-aware Echo State Network (ESN) infers the parameterized dynamics. Second, we derive the adjoint of the ESN to compute the sensitivity of time-averaged cost functionals. Third, we propose the Thermoacoustic Echo State Network (T-ESN), which hard constrains the physical knowledge in the network architecture. Fourth, we apply the framework to a variety of nonlinear thermoacoustic regimes of a prototypical system. We show that the T-ESN accurately infers the correct adjoint sensitivities of the time-average acoustic energy with respect to the flame parameters. The results are robust to noisy data, from periodic, through quasiperiodic, to chaotic regimes. The inferred adjoint sensitivities are employed to suppress an instability via steepest descent. We show that a single network predicts the nonlinear bifurcations on unseen scenarios, which allows it to converge to the minimum of the acoustic energy. This work opens new possibilities for gradient-based data-driven design optimization.
@@ I Introduction
Engineers and scientists strive to design flow and acoustic systems that behave in an optimal way [e.g., 1, 2, 3, 4, 5]. Finding the global optimum is challenging because the temporal dynamics can be nonlinear, time-delayed, and chaotic, which makes the objective function non-convex [e.g, 6, 7]. A practical approach is to solve the optimization problem by gradient-descent, in which the objective function is optimized along the gradient direction until a local optimum is found. The gradient (also known as sensitivity) offers the information about how the quantity of interest changes with the system’s parameters. Central to gradient-based optimization methods are adjoint methods. Adjoint methods enable the accurate computation of the gradient with a computational cost that is not affected by the number of design parameters. For example, if the system has $N$ parameters, adjoint methods compute the gradient of a objective function with only one calculation. This is is contrast with tangent linear or finite difference methods, which would require $N$ computations [e.g., 8, 9, 10, 2, 3, 5].
@@
Although adjoint methods are excellent tools, they are problem dependent, i.e., they depend on the system’s equations. First, the adjoint sensitivity is only as accurate as the model assumptions. Low-order models are computationally cheap to evaluate, however the uncertainties in the model can lead to errors in the sensitivity computations, especially in nonlinear systems that are sensitive to perturbations. Second, the system must be linearized, i.e., by computing the Jacobian. This means the computation of the adjoint solver needs to be performed for different fluid problems, numerical schemes, and systems. Automatic differentiation [16, 17] and surrogate models  [18] help the automation of the process, but they still require a code that may be problem dependent.
@@
To bypass the need of system’s equations, a data-driven model can be employed. The key idea of this paper is to learn an approximation of the dynamical systems from observables, and then compute the adjoint of it. Recurrent neural networks are nonlinear dynamical systems, which infer the governing dynamics from observables (data). We employ a type of reservoir computer, the Echo State Network (ESN) [19], which is a universal approximator of the dynamics [20]. The ESN benefits from straightforward training because it does not require backpropagation. Its Jacobian can also be computed in a straightforward way due to the shallow architecture [21]. ESNs have been shown to successfully (i) make time-accurate short-term forecasts [22, 23, 24], (ii) reconstruct unmeasured variables [25, 26], (iii) learn ergodic properties, such as Lyapunov exponents, and long-term statistics of chaotic flows [27, 7, 21], and (iv) infer model errors in thermoacoustics [28]. In recent advances, low-order representations of turbulence were learned from high-dimensional data, and the dynamics of this low-dimensional system was propagated with the ESN [29]. In this paper, we take an additional step. Because in order to perform design optimization of system’s parameters, we require a parameter-aware extension of the Echo State Network [30], which learns the parameterized dynamics and can operate across parameter regimes.
@@
We apply the data-driven sensitivity framework to nonlinear thermoacoustic oscillations. These are phenomena that can be detrimental in the manufacturing of gas turbines [31], and rocket engines [32]. In thermoacoustic systems, if the heat release and acoustic pressure are sufficiently in phase, self-sustained oscillations can occur [33, 34]. Therefore, their prediction, control, and suppression are important both during design and during operation. This is especially challenging because thermoacoustics exhibit nonlinear behaviour and bifurcations to periodic, quasiperiodic, and chaotic regimes [35], which motivates the development of methods for nonlinear analysis. Preliminary design of thermoacoustic systems is performed by linear stability [31, 36]. Linear stability analysis determines the growth rates of infinitesimal perturbations to the mean flow, i.e., the eigenvalues of the linearized system of equations around the mean flow [37]. To name a few examples, adjoint-based linear stability analysis enabled calculation of sensitivity to parameters and passive control mechanisms [38], optimal placement and tuning of acoustic dampers [39] and stabilization of thermoacoustic oscillations by geometry changes [40]. An in-depth review of development of adjoint methods for thermoacoustics can be found in [5, 37]. Beyond linear stability analysis, adjoint methods were used for weakly nonlinear analysis to predict subcritical Hopf bifurcations [41], and Floquet analysis to calculate the stability and the period of limit cycle oscillations [5]. The stability of chaotic thermoacoustics was studied by Huhn and Magri [6] using covariant Lyapunov analysis, who analytically showed that this is equivalent to eigenvalue and Floquet analysis for fixed point and limit cycle solutions, respectively. As a result, Huhn and Magri [6] achieved gradient-based minimization of acoustic energy in chaotic thermoacoustics using shadowing methods. The application of this method was challenged by the discontinuities observed in the acoustic energy when bifurcations occur, computational cost of shadowing methods, and non-hyperbolicity of thermoacoustics for some regimes. To circumvent these issues, a gradient-free optimization method based on Bayesian optimization was proposed [7]. Among applications of data-driven methods in thermoacoustics, the ESN was previously used to predict the nonlinear dynamics [7] and calculate the Lyapunov exponents [42] of a chaotic regime, and infer the model error of a low-order model [28].
@@
The overarching objective of this paper is to address the challenge of developing code-specific adjoint solvers, which requires the system’s equations and the Jacobian to be known. We propose data-driven nonlinear direct-adjoint looping to compute parametric sensitivities, and apply this to a prototypical thermoacoustic system. The specific objectives are to build a robust parameter-aware ESN architecture for this purpose, derive its adjoint, and demonstrate the data-driven inference of parameterized dynamics and associated adjoint sensitivities. We refer to this architecture as the thermoacoustic ESN and show that (i) it learns the parameterized dynamics of nonlinear, time-delayed thermoacoustics, (ii) predicts bifurcations, (iii) accurately obtains the adjoint sensitivities, (iv) the acoustic energy is minimized in a gradient-based optimization framework, and (v) demanding scenarios such as training with noisy or chaotic data can be handled.
@@
This paper is organized as follows. In Sec. II, we give a background on the mathematical framework for sensitivity analysis and the parameter-aware ESN. In Sec. III, we derive the adjoint of the ESN. In Sec. IV, we present the nonlinear and time-delayed thermoacoustic model, namely the Rijke tube. In Sec. V, we introduce the time-delayed ESN designed for thermoacoustics and derive its adjoint. In Sec. VI, we show the performance of the ESN in terms of prediction and accuracy of the learned adjoint sensitivities, and employ it for the gradient-based design optimization of thermoacoustics in Sec. VII. We address training with noisy data in Sec. VIII, and learning chaotic regimes in Sec. IX. Section X concludes the paper.
@@ II.1 Sensitivity Analysis
where $\mathbf{x}\in\mathbb{R}^{n_{x}}$ is the state vector, $\dot{\mathbf{x}}$ is its time-derivative, $\mathbf{p}\in\mathbb{R}^{n_{p}}$ is the parameters’ vector, and $\mathbf{F}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{x}}$ is the operator that describes the system’s evolution. We assume that $\mathbf{F}$ is smooth so that the Jacobian of the system, i.e., $\partial\mathbf{F}/\partial\mathbf{x}$, is defined everywhere, which is a requirement for gradient-based sensitivity analysis. We wish to optimize an objective functional, $\mathcal{J}$, which is a time-average of $\tilde{\mathcal{J}}$ [5]
@@
$\dot{\mathbf{q}}^{+}$ $=-\frac{\partial\tilde{\mathcal{J}}}{\partial\mathbf{x}}^{\top}-\frac{\partial\mathbf{F}}{\partial\mathbf{x}}^{\top}\mathbf{q},$ (5b)
@@ II.2 Parameter-aware Echo State Networks
Given the evolution equations of a dynamical system (1), we can compute its adjoint (5). Developing the adjoint solver however requires the system Jacobian, $\partial\mathbf{F}/\partial\mathbf{x}$. This Jacobian is code specific. The key idea of this paper is to eliminate code-specific adjoint solvers by having a machine that learns the dynamics from data, and is itself a dynamical system, the Jacobian of which is readily available independent of the system in question. Recurrent neural networks (RNNs) are well-suited for this task, as they are neural network architectures that can process sequential data and make forecasts of timeseries from dynamical systems. Although the dynamical systems we consider (1) are continuous in time, RNNs accept data as a sequence in discrete time. RNNs possess memory through a hidden state and thus they can learn temporal relationships. The hidden state enables the RNNs to be framed as discrete dynamical systems [21]. In this work, we utilize the Echo State Network (ESN) [43], which is a type of RNN that is straightforward to train because it does not require backpropagation during training. The ESN is a universal approximator [20] and it follows this dynamic equation
@@
where $\mathbf{y}\in\mathbb{R}^{n_{y}}$ is the target output vector, the dynamics of which we aim to model, $\mathbf{\hat{y}}\in\mathbb{R}^{n_{y}}$ is the predicted output vector, and $\mathbf{W}_{out}\in\mathbb{R}^{n_{y}\times n_{r}}$ is the output matrix that maps the reservoir state to the output vector. The matrices $\mathbf{W}_{in}$ and $\mathbf{W}$ are sparse, randomly generated, and not trained, whilst $\mathbf{W}_{out}$ is trained via ridge regression [19]. The input matrix is scaled as $\mathbf{W}_{in}=\sigma_{in}\tilde{\mathbf{W}}_{in}$, where $\tilde{\mathbf{W}}_{in}$ is constructed such that each row contains only one nonzero element drawn from a uniform distribution $\sim\mathcal{U}(-1,1)$, i.e., the inputs are activated individually. The state matrix $\mathbf{W}$ describes the reservoir connections, and is scaled as $\mathbf{W}=\rho\tilde{\mathbf{W}}$, where $\tilde{\mathbf{W}}$ is an Erdős-Renyi matrix with a given average number of nonzero elements per row (connectivity) drawn from a uniform distribution $\sim\mathcal{U}(-1,1)$. The scalar $\rho$ is the spectral radius of $\mathbf{W}$. In this architecture, the reservoir state variables “echo” the input, and so the reservoir holds memory. During training $\mathbf{y}_{in}$ is fed into the network in open-loop and the evolution of the reservoir state is saved. The reservoir state is initialized to a zero vector. An initial time window is used for the washout stage, which eliminates the effects of the initial condition and is discarded in training. The training is performed by solving for $\mathbf{W}_{out}$ [19]
@@
where $\mathbf{R}=[\mathbf{r}(1)\;\mathbf{r}(2)\;\dots\;\mathbf{r}(N_{t})],\;\mathbf{R}\in\mathbb{R}^{n_{r}\times N_{t}}$, $\mathbf{Y}=[\mathbf{y}(1)\;\mathbf{y}(2)\;\dots\;\mathbf{y}(N_{t})],\;\mathbf{Y}\in\mathbb{R}^{n_{y}\times N_{t}}$, and $\mathbf{I}$ is the identity matrix. Tuning an ESN involves searching for optimal hyperparameters that consist of input matrix scaling $\sigma_{in}$, the spectral radius of the state matrix $\rho$ (typically $\leq 1$ to satisfy the echo state property), the leak rate $\alpha$, and the Tikhonov regularizer $\lambda$. These hyperparameters are determined by validation. A detailed explanation of training and robust validation of Echo State Networks can be found in [24]. After training, the ESN can run autonomously in closed-loop by setting $\mathbf{y}_{in}(i)=\mathbf{\hat{y}}(i)$.
@@
where (;)(;)( ; ) denotes vertical concatenation, $\mathbf{p}\in\mathbb{R}^{n_{p}}$ denotes the system’s parameters and are fed to the network through the input channel, which we assume to be constant in time for each system. The input matrix $\mathbf{W}_{in}=[\mathbf{W}_{in}^{y}\;\mathbf{W}_{in}^{p}],\;\mathbf{W}_{in}^{y}\in\mathbb{R}^{n_{r}\times n_{y}},\;\mathbf{W}_{in}^{p}\in\mathbb{R}^{n_{r}\times n_{p}}$ is constructed such that the parameters are fully connected with all the reservoir nodes, i.e., the weights associated with the parameters $\mathbf{W}_{in}^{p}$ are nonzero for all rows. The parameter-aware ESN adds two additional hyperparameters per physical parameter $p_{i}$. A scalar $k_{p_{i}}$ shifts the parameter $p_{i}$ as $p_{i}-k_{p_{i}}$, and a scalar $\sigma_{p_{i}}$ scales it as $\sigma_{p_{i}}(p_{i}-k_{p_{i}})$. Hence, we apply the input scaling $\sigma_{in}$ on $\mathbf{W}_{in}^{y}$ only. The training of the parameter-aware ESN is performed as the standard ESN (8). The dataset contains data from different parameter regimes, which means that we need to save the evolution of the reservoir state for each training regime, and concatenate as $\mathbf{R}=[\mathbf{R}^{1}\;\mathbf{R}^{2}\;\dots,\mathbf{R}^{N_{tr}}]$ (likewise, the outputs are concatenated as $\mathbf{Y}=[\mathbf{Y}^{1}\;\mathbf{Y}^{2}\;\dots\mathbf{Y}^{N_{tr}}]$, where $N_{tr}$ is the number of training regimes.)

### v2

@@ Abstract
Adjoint methods have been the pillar of gradient-based optimization for decades. They enable the accurate computation of a gradient (sensitivity) of a quantity of interest with respect to all system’s parameters in one calculation. When the gradient is embedded in an optimization routine, the quantity of interest can be optimized for the system to have the desired behaviour. Adjoint methods, however, require the system’s governing equations and their Jacobian. In this paper, we propose a computational strategy to infer the adjoint sensitivities from data when the governing equations might be unknown (or partly unknown), and noise might be present. The key component of this strategy is an echo state network, which learns the dynamics of nonlinear regimes with varying parameters, and it evolves dynamically via a hidden state. Although the framework does not make assumptions on the dynamical system, we focus on thermoacoustics, which are governed by nonlinear and time-delayed systems. First, we show that a parameter-aware Echo State Network (ESN) infers the parameterized dynamics. Second, we derive the adjoint of the ESN to compute two types of sensitivity: (i) parameter sensitivity, which is the gradient of a time-averaged cost functional with respect to physical/design parameters of the system; and (ii) initial condition sensitivity, which is the gradient of a cost functional of the final state with respect to the initial condition. Third, we propose the Thermoacoustic Echo State Network (T-ESN), which embeds the physical knowledge in the network architecture for improved generalization. Fourth, we apply the framework to a variety of nonlinear thermoacoustic regimes of a prototypical system. We show that the T-ESN accurately infers the correct adjoint sensitivities of the acoustic energy with respect to the flame parameters and initial conditions. The results are robust to noisy data, from periodic, through quasiperiodic, to chaotic regimes. The inferred adjoint sensitivities are employed to suppress an instability via steepest descent. We show that a single network predicts the nonlinear bifurcations on unseen scenarios, which allows it to converge to the minimum of the acoustic energy. This work opens new possibilities for gradient-based data-driven design optimization.
@@ I Introduction
Engineers and scientists strive to design flow and acoustic systems that behave in an optimal way [e.g., 1, 2, 3, 4, 5]. Finding the global optimum is challenging because the temporal dynamics can be nonlinear, time-delayed, and chaotic, which makes the objective function non-convex [e.g, 6, 7]. A practical approach is to solve the optimization problem by gradient-descent, in which the objective function is optimized along the gradient direction until a local optimum is found. The gradient (also known as sensitivity) offers the information about how the quantity of interest changes with the system’s parameters. Central to gradient-based optimization methods are adjoint methods. Adjoint methods enable the accurate computation of the gradient with a computational cost that is not affected by the number of design parameters. For example, if the system has $N$ parameters, adjoint methods compute the gradient of a objective function with only one calculation. This is in contrast with tangent linear or finite difference methods, which would require $N$ computations [e.g., 8, 9, 10, 2, 3, 5].
@@
To bypass the need of the system’s equations, a data-driven model can be employed. The key idea of this paper is to learn an approximation of the dynamical systems from observables, and then compute the adjoint of it. This approach is useful when working with experimental data, or data whose equations are not fully known. To infer the governing dynamics from observables (data), we employ a type of reservoir computer, the Echo State Network (ESN) [16], which is a universal approximator and a type of recurrent neural network (RNN) [17]. The ESN’s training is straightforward because it does not require backpropagation. Its Jacobian can also be computed in a straightforward way due to the shallow architecture [18]. ESNs have been shown to successfully (i) make time-accurate short-term forecasts [19, 20, 21], (ii) reconstruct unmeasured variables [22, 23], (iii) learn ergodic properties, such as Lyapunov exponents, and long-term statistics of chaotic flows [24, 7, 18], and (iv) infer model errors in thermoacoustics [25]. In recent advances, low-order representations of turbulence were learned from high-dimensional data, and the dynamics of this low-dimensional system was propagated with the ESN [26]. In this paper, we take an additional step. Because in order to perform design optimization of system’s parameters, we require a parameter-aware extension of the Echo State Network [27], which learns the parameterized dynamics and can operate across parameter regimes. In a recent study, [28] demonstrated the proposed data-driven sensitivity framework on a prototypical chaotic system, and obtained the adjoint sensitivities of long-time averages to the system’s parameters.
@@
In this paper, we develop a data-driven sensitivity framework, with an application on nonlinear and time-delayed thermoacoustic oscillations. These are phenomena that can be detrimental in the manufacturing of gas turbines [29], and rocket engines [30]. In thermoacoustic systems, if the heat release and acoustic pressure are sufficiently in phase, self-sustained oscillations can occur [31, 32]. Therefore, their prediction, control, and suppression are important both during design and during operation. This is especially challenging because thermoacoustics exhibit nonlinear behaviour and bifurcations to periodic, quasiperiodic, and chaotic regimes [33], which motivates the development of methods for nonlinear analysis. Preliminary design of thermoacoustic systems is performed by linear stability [29, 34]. Linear stability analysis determines the growth rates of infinitesimal perturbations to the mean flow, i.e., the eigenvalues of the linearized system of equations around the mean flow [35]. To name a few examples, adjoint-based linear stability analysis enabled calculation of sensitivity to parameters and passive control mechanisms [36], optimal placement and tuning of acoustic dampers [37] and stabilization of thermoacoustic oscillations by geometry changes [38]. An in-depth review of development of adjoint methods for thermoacoustics can be found in [5, 35]. Beyond linear stability analysis, adjoint methods were used for weakly nonlinear analysis to predict subcritical Hopf bifurcations [39], and Floquet analysis to calculate the stability and the period of limit cycle oscillations [5]. The stability of chaotic thermoacoustics was studied by [6] using covariant Lyapunov analysis, who analytically showed that this is equivalent to eigenvalue and Floquet analysis for fixed point and limit cycle solutions, respectively. As a result, [6] achieved gradient-based minimization of acoustic energy in chaotic thermoacoustics using shadowing methods. The application of this method was challenged by the discontinuities observed in the acoustic energy when bifurcations occur, computational cost of shadowing methods, and non-hyperbolicity of thermoacoustics for some regimes. To circumvent these issues, a gradient-free optimization method based on Bayesian optimization was proposed [7]. Among applications of data-driven methods in thermoacoustics, the ESN was previously used to predict the nonlinear dynamics [7] and calculate the Lyapunov exponents [40] of a chaotic regime, and infer the model error of a low-order model [25].
@@
The overarching objective of this paper is to address the challenge of developing code-specific adjoint solvers, which requires the system’s equations and the Jacobian to be known. We propose data-driven nonlinear direct-adjoint looping to compute sensitivities to the system’s parameters and initial conditions, and apply this to a prototypical thermoacoustic system. The specific objectives are to introduce a robust parameter-aware ESN architecture for this purpose, derive its adjoint, and demonstrate the data-driven inference of parameterized dynamics and associated adjoint sensitivities. We refer to this architecture as the thermoacoustic ESN and show that (i) it learns the parameterized dynamics of nonlinear, time-delayed thermoacoustics, (ii) predicts bifurcations, (iii) accurately obtains the adjoint sensitivities, (iv) the acoustic energy is minimized in a gradient-based optimization framework, and (v) demanding scenarios such as training with noisy or chaotic data can be handled.
@@
This paper is organized as follows. In Sec. II, we give a background on the mathematical framework for sensitivity analysis and the parameter-aware ESN. In Sec. III, we derive the adjoint of the ESN. In Sec. IV, we present the nonlinear and time-delayed thermoacoustic model, namely the Rijke tube. In Sec. V, we introduce the time-delayed ESN designed for thermoacoustics and derive its adjoint. In Sec. VI, we show the performance of the ESN in terms of prediction and accuracy of the learned adjoint sensitivities, and employ it for the gradient-based design optimization of thermoacoustics in Sec. VII. We address training with noisy data in Sec. VIII, and learning chaotic regimes in Sec. IX. We show the sensitivity to initial conditions in Section X. A discussion is offered in Section XI. Section XII concludes the paper.
@@ II.1 Sensitivity Analysis
where $\mathbf{x}\in\mathbb{R}^{n_{x}}$ is the state vector, $\dot{\mathbf{x}}$ is its time derivative, $\mathbf{p}\in\mathbb{R}^{n_{p}}$ is the parameters’ vector, and $\mathbf{F}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{x}}$ is the operator that describes the system’s evolution. We assume that $\mathbf{F}$ is smooth so that the Jacobian of the system, i.e., $\partial\mathbf{F}/\partial\mathbf{x}$, is defined everywhere, which is a requirement for gradient-based sensitivity analysis. We wish to optimize an objective functional, $\mathcal{J}$, which is a time-average of $\tilde{\mathcal{J}}$ [5]
@@
$\dot{\mathbf{q}}^{+}$ $=-\frac{\partial\tilde{\mathcal{J}}}{\partial\mathbf{x}}^{\top}-\frac{\partial\mathbf{F}}{\partial\mathbf{x}}^{\top}{\color[rgb]{0,0,0}\mathbf{q}^{+}},$ (5b)
@@ II.2 Parameter-aware Echo State Networks
Given the evolution equations of a dynamical system (1), we can compute its adjoint (5). Developing the adjoint solver however requires the system Jacobian, $\partial\mathbf{F}/\partial\mathbf{x}$. This Jacobian is code specific. The key idea of this paper is to eliminate code-specific adjoint solvers by having a machine that learns the dynamics from data, and is itself a dynamical system, the Jacobian of which is readily available independent of the system under investigation. Recurrent neural networks (RNNs) are well-suited for this task, as they are neural network architectures that can process sequential data and make forecasts of timeseries from dynamical systems. Although the dynamical systems we consider in Eq. (1) are continuous in time, RNNs accept data as a sequence in discrete time. RNNs possess memory through a hidden state and thus they can learn temporal relationships. The hidden state enables the RNNs to be framed as discrete dynamical systems [18]. In this work, we utilize the Echo State Network (ESN) [41], which is a type of RNN that is straightforward to train because it does not require backpropagation during training. The ESN is a universal approximator [17] and it follows this dynamic equation
@@
where $\mathbf{y}\in\mathbb{R}^{n_{y}}$ is the target output vector, the dynamics of which we aim to model, $\mathbf{\hat{y}}\in\mathbb{R}^{n_{y}}$ is the predicted output vector, and $\mathbf{W}_{out}\in\mathbb{R}^{n_{y}\times n_{r}}$ is the output matrix that maps the reservoir state to the output vector. The matrices $\mathbf{W}_{in}$ and $\mathbf{W}$ are sparse, randomly generated, and not trained, whilst $\mathbf{W}_{out}$ is trained via ridge regression [16]. The input matrix is scaled as $\mathbf{W}_{in}=\sigma_{in}\tilde{\mathbf{W}}_{in}$, where $\tilde{\mathbf{W}}_{in}$ is constructed such that each row contains only one nonzero element drawn from a uniform distribution $\sim\mathcal{U}(-1,1)$, i.e., the inputs are activated individually. The state matrix $\mathbf{W}$ describes the reservoir connections, and is scaled as $\mathbf{W}=\rho\tilde{\mathbf{W}}$, where $\tilde{\mathbf{W}}$ is an Erdős-Renyi matrix with a given average number of nonzero elements per row (connectivity) drawn from a uniform distribution $\sim\mathcal{U}(-1,1)$, which is scaled to have a unity spectral radius. So, the scalar $\rho$ is set as the spectral radius of $\mathbf{W}$. In this architecture, the reservoir state variables “echo” the input, and so the reservoir holds memory. During training $\mathbf{y}_{in}$ is fed into the network in open-loop and the evolution of the reservoir state is saved. The reservoir state is initialized to a zero vector. An initial time window is used for the washout stage, which eliminates the effects of the initial condition and is discarded in training. The training is performed by solving for $\mathbf{W}_{out}$ [16]
@@
where $\mathbf{R}=[\mathbf{r}(1)\;\mathbf{r}(2)\;\dots\;\mathbf{r}(N_{t})],\;\mathbf{R}\in\mathbb{R}^{n_{r}\times N_{t}}$, $\mathbf{Y}=[\mathbf{y}(1)\;\mathbf{y}(2)\;\dots\;\mathbf{y}(N_{t})],\;\mathbf{Y}\in\mathbb{R}^{n_{y}\times N_{t}}$, and $\mathbf{I}$ is the identity matrix. Tuning an ESN involves searching for optimal hyperparameters that consist of input matrix scaling $\sigma_{in}$, the spectral radius of the state matrix $\rho$ (typically $\leq 1$ to satisfy the echo state property), the leak rate $\alpha$, and the Tikhonov coefficient $\lambda$. These hyperparameters are determined by validation. A detailed explanation of training and robust validation of Echo State Networks can be found in [21]. After training, the ESN can run autonomously in closed-loop by setting $\mathbf{y}_{in}(i)=\mathbf{\hat{y}}(i)$.
@@
where (;)(;)( ; ) denotes vertical concatenation, $\mathbf{p}\in\mathbb{R}^{n_{p}}$ denotes the system’s parameters and are fed to the network through the input channel, which we assume to be constant in time for each system. The input matrix $\mathbf{W}_{in}=[\mathbf{W}_{in}^{y}\;\mathbf{W}_{in}^{p}],\;\mathbf{W}_{in}^{y}\in\mathbb{R}^{n_{r}\times n_{y}},\;\mathbf{W}_{in}^{p}\in\mathbb{R}^{n_{r}\times n_{p}}$ is constructed such that the parameters are fully connected with all the reservoir nodes, i.e., the weights associated with the parameters $\mathbf{W}_{in}^{p}$ are nonzero for all rows. The parameter-aware ESN adds two additional hyperparameters per physical parameter $p_{i}$. A scalar $k_{p_{i}}$ shifts the parameter $p_{i}$ as $p_{i}-k_{p_{i}}$, and a scalar $\sigma_{p_{i}}$ scales it as $\sigma_{p_{i}}(p_{i}-k_{p_{i}})$. Hence, we apply the input scaling $\sigma_{in}$ on $\mathbf{W}_{in}^{y}$ only. The training of the parameter-aware ESN is performed as the standard ESN (8). The dataset contains data from different parameter regimes, which means that we need to save the evolution of the reservoir state for each training regime, and concatenate as $\mathbf{R}=[\mathbf{R}^{1}\;\mathbf{R}^{2}\;\dots,\mathbf{R}^{N_{tr}}]$ (likewise, the outputs are concatenated as $\mathbf{Y}=[\mathbf{Y}^{1}\;\mathbf{Y}^{2}\;\dots\mathbf{Y}^{N_{tr}}]$), where $N_{tr}$ is the number of training regimes.